OpenStudy (anonymous):

Given the two functions below, what is the x-intercept of g(x)? f(x) = x2 g(x) = f(x –7)

OpenStudy (yrelhan4):

what will be gx ?

OpenStudy (anonymous):

If your asking me i don't know. lol That's the whole question

OpenStudy (yrelhan4):

fx = x^2 right?

OpenStudy (anonymous):

Yes

OpenStudy (yrelhan4):

so.. g(x)= f(x-7) .. that basically means g(x) is equal to f(y) where value of y would be x-7.. so all you need to do is..put x-7 in place of x in f(x).. that would give you g(x)=(x-7)^2.. yeah?

OpenStudy (irishboy123):

f(g) = (x-7)^2 y=0 @ x=7

OpenStudy (yrelhan4):

wrong anu.. put g(x)=0 and find the value of x..

OpenStudy (anonymous):

Yes @yrelhan4 u r right its a typo :)

OpenStudy (anonymous):

Yes your right @yrelhan4

OpenStudy (anonymous):

thanks everyone :)